If 5 grams of nitrogen is oxidized in air, then what is the empirical formula of compound, if it's mass is 19.3 grams?
1
Expert's answer
2013-01-08T12:03:28-0500
General equation for this process is next:

N2 + xO2 = 2NOx

let's find x

5 19.3
N2 + xO2 = 2NOx
28 2(14+16x)

2(14 + 16x)*5 = 28* 19.3
2(14 + 16x)*5 = 540.4
14 + 16x = 54.04
16x = 40
x = 2.5

so empiric formula is NO2.5, or N2O5